Weber Interstate Paving Co. had $450 million of sales and $225 million of fixed assets last year, so its FA/Sales ratio was 50%. However, its fixed assets were used at only 85% of capacity. If the company had been able to sell off enough of its fixed assets at book value so that it was operating at full capacity, with sales held constant at $450 million, how much cash (in millions) would it have generated?

Select the correct answer.

a. $35.45
b. $38.85
c. $33.75
d. $30.35
e. $37.15
Answer #1

At 85% capacity, sales of 450 million
At 100% capacity, assets required for the same sales=225*0.85=191.25

Cash generated=225-191.25=33.75